Here is a problem I have put through a few times, but getting difference answers and a bit confused. Hope someone can show me how to work it out.
I need to solve for x and show proof:
-4 (2x -3) = 8x + 5
Thanks,
Sher

Answer by stanbon(48535)

-4 (2x -3) = 8x + 5
Use distribution to get:
-8x+12 = 8x + 5
Add 8x and subtract 5 from both sides to get:
16x=7
Divide both sides by 16 to get:
x=7/16
Proof:
-4(2(7/16)-3) = 8(7/16) + 5
-4(7/8-24/8) =7/2+10/2
-4(-17/8) = 17/2
17/2= 17/2
Cheers,
Stan H.